OpenStudy (anonymous):

If the product of 14 and a number is decreased by 7, and then the result is divided by 5, the answer is 7. What is the unknown number? Select the best answer from the choices provided. 0 2 3 6

OpenStudy (anonymous):

answer=(3) (14x-7)/5=7 cross multiply 14x-7=7*5 14x-7=35 14x=35+7 14x=42 x=42/14 x=3 ans. Proof: (14*3-7)/5=7 (42-7)/5=7 35=7*5 35=35